OpenStudy (anonymous):

. At a certain time of day, a 6 ft man casts a 4 ft shadow. At the same time of day, how tall is a tree that casts an 18 ft shadow?

OpenStudy (anonymous):

Let x be the height of the tree. Make a proportion. x/ 18 = 6/4 6/4 * 18 =27